The sum of Andy and Mandy's age is 20. The product of their age is 96. If Andy is older then Mandy how old is andy use equations with a letter representing the unknown.
Answer by alexlove054@gmail.com(10)

there's a mistake in the question the sum of their age is supposed to 20.
let their ages be x and y
x y=20.....1
xy=96.......2
make x the subject of the formula in equation 1
x=20-y......3
put in equation 2
(20-y)y=96
20y-y^2=96
take the left side to the right
y^2-20 96=0
factorize
y^2-12y-8y 96=0
y(y-12)-8(y-12)=0
y=8 and y=12
put y in equation 3
x=20-8=12
x=12
and y =8